# Diffusion approximation of the stationary distribution of a two-level single server queue

### v7

@@ 3.2 Main results
We are now ready to present a main result, which will be proved in Section 5. Recall that $L^{(r)}$ is a random variable subject to the stationary distribution of the $r$-th 2-level $GI/G/1$ queue. Define probability distribution $\nu^{(r)}$ under the assumptions (3.a)–(3.e) as
@@
We refer to this $\nu^{(r)}$ as a scaled stationary distribution. We decompose $\nu^{(r)}$ using finite measures $\nu^{(r)}_{i}$ and conditional probability distributions $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$, which are defined as, for $B\in{\mathcal{B}}(\mathbb{R}_{+})$,
$\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;\ell^{(r)}_{0}<L^{(r)}),$
$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|\ell^{(r)}_{0}<L^{(r)}).$
@@
$h_{1|cd}(x)=1(0\leq x\leq\ell_{0})\times\left\{\begin{array}[]{ll}\displaystyle 1/\ell_{0},&b_{1}=0,\\ \displaystyle\frac{e^{\beta_{1}\ell_{0}}\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}e^{-\beta_{1}x},&b_{1}\not=0,\end{array}\right.$ (3.4)
$h_{2|cd}(x)=1(x>\ell_{0})\beta_{2}e^{-\beta_{2}(x-\ell_{0})}.$ (3.5)
@@
    $\widetilde{A}_{1}=\begin{cases}\displaystyle\frac{2b_{2}\ell_{0}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{0}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)},\;&b_{1}\not=0,\end{cases}\qquad\widetilde{A}_{2}=1-\widetilde{A}_{1}.$ (3.10)
@@
Hence, by Little’s law, we have $\rho^{(r)}=\mathbb{P}(L^{(r)}>0)=1-\mathbb{P}(L^{(r)}=0)$. In diffusion applications of a single server queue, the mean queue length is estimated through $(1-\rho^{(r)})\mathbb{E}[L^{(r)}]$ as $r\downarrow 0$. The following corollary computes this limit, whose details are given in Appendix A.1.
@@
=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{0}(b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{0}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\vspace{1ex}\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{0}}-1)+2b_{1}b_{2}\ell_{0}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.15)
@@
$\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{\beta_{2}\ell_{0}}{1+\beta_{2}\ell_{0}},&b_{1}=0,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.16)
@@
Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed. However, for a double check, we give its direct proof in Appendix A.2, which is helpful to convince (ii) of Theorem 3.1. This proof is challenging because the exponential model requires a Markov chain with two-dimensional state space to correctly describe the case that the new arrival will find a system with less than $\ell_{0}$ after multiple consecutive departures when the previous arrival found a queue length of greater than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1). This Markov chain will be detailed in Appendix A.2.
@@ 4 BAR’s for the 2-level $GI/G/1$ queue
In this section, we prepare auxiliary notations and results for proving Theorem 3.1. A major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$. Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions, which will be defined below. This method is called a BAR approach, and has been used in [3, 4, 10]. Recently, this approach is further studied in [12].
Throughout this section, we assume (3.a)–(3.e), but (3.f) is not assumed. The proof of Theorem 3.1 will be done in Section 5.
@@ 4.1 Palm distribution and BAR
Then, for each $r\in(0,1]$, $X^{(r)}(\cdot)$ and the increments of $N^{(r)}_{e}$ and $N^{(r)}_{s}$ are jointly stationary. Since $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite by Lemma 3.1, we can define probability distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ on $(\Omega,{\mathcal{F}})$ by
@@
We now derive a BAR for $X^{(r)}(\cdot)$ using the framework proposed in [12]. For this, we introduce notations for describing the dynamics of $X^{(r)}(\cdot)$. Let $S$ be its state space, that is, $S=\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded and continuous functions from $S$ to $\mathbb{R}_{+}$ which are partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}$ is partial derivative from the right. Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
${\mathcal{H}}f(z,{\boldsymbol{x}})=-\frac{\partial^{+}}{\partial x_{2}}f(z,{\boldsymbol{x}})-(c_{1}1(1\leq z\leq\ell_{0})+c_{2}1(z>\ell_{0}))\frac{\partial^{+}}{\partial x_{2}}f(z,{\boldsymbol{x}}),\quad(z,{\boldsymbol{x}})\in S,$
where ${\boldsymbol{x}}=(x_{1},x_{2})\in\mathbb{R}_{+}^{2}$.
@@
$X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N^{(r)}_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{0},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{0},\end{cases}$
@@
where $n_{e,1}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)<\ell^{(r)}_{0})$ and $n_{e,2}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)\geq\ell^{(r)}_{0})$. We call $X^{(r)}_{1}(t)$ an intermediate state. Note that $X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N^{(r)}_{s}(t)))$ when $\Delta N^{(r)}_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N^{(r)}_{s}(t)=0$.
@@
$\Delta_{e}f(X^{(r)}(t))=f(X^{(r)}_{1}(t))-f(X^{(r)}(t-)),$
$\Delta_{s}f(X^{(r)}(t))=f(X^{(r)}(t))-f(X^{(r)}_{1}(t)).$
Then, from the elementary observation of the time evolution of $X^{(r)}(t)$, we have
$f(X^{(r)}(t)=f(X^{(r)}(0))+\int_{0}^{t}{\mathcal{H}}f(X^{(r)}(u))du+\sum_{u\in(0,t]}[\Delta_{e}f(X^{(r)}(u))+\Delta_{s}f(X^{(r)}(u))],$
where the summation is well defined because $\Delta_{e}f(X^{(r)}(u))$ and $\Delta_{s}f(X^{(r)}(u))$ vanish except for finitely many $u\in(0,t]$. Take the expectation of the above equation for $t=1$ by $\mathbb{P}$, then the stationarity of $X(\cdot)$ and the definition of the Palm distribution yield
@@
$\mathbb{E}\left[\sum_{u\in(0,1]}\Delta_{v}f(X^{(r)}(u))]\right]=\alpha^{(r)}_{v}\mathbb{E}^{(r)}_{v}\left[\Delta_{v}f(X^{(r)}(0))\right],\qquad v=e,s.$
@@
$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]=0,\qquad f\in C_{b}^{p},$ (4.2)
The formula (4.1) is a special case of the rate conservation law (e.g., see [9]), and referred as a basic adjoint relationship, BAR in short, then we call analysis based on this BAR as a BAR approach. A major difficulty in this BAR approach is to evaluate the expectations under the Palm distributions (the last two terms in (4.1)). This evaluation will be done for the 2-level $GI/G/1$ queue in Section 4.3.
@@
In using Palm distributions, we must be careful about the following two points.
  • (i)
    $X^{(r)}(0-)$ is not defined under $\mathbb{P}$, but its distribution is well defined under the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{e}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$. So, $X^{(r)}(0-)$ is a random variable well defined under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$. Hence, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined under the Palm distributions. If we define the stationary process $X^{(r)}(\cdot)$ on the time axis (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ are well defined even under $\mathbb{P}$.
  • (ii)
    Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N^{(r)}_{e}$ is counted at time 00 and just before when $N^{(r)}_{s}$ is counted at time 00. This means that we take a so called arrival first framework. As for the intermediate state, see [12] for its details. We use $L^{(r)}(0-)$ under $\mathbb{P}^{(r)}_{e}$, while $L^{(r)}(0)$ under $\mathbb{P}^{(r)}_{s}$. This prevents to directly refer to the intermediate state $X^{(r)}_{1}(0)$ in our arguments. Of course, the intermediate state is not needed if $N^{(r)}_{e}$ and $N^{(r)}_{s}$ do not have a simultaneous count, but such restriction is avoided by introducing $X^{(r)}_{1}(0)$.
@@
Under conditions (3.a) and (3.c), $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$, and
@@
$\alpha^{(r)}_{e}=\frac{\lambda^{(r)}_{1}\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}\mathbb{P}^{(r)}_{e}(L^{(r)}(0-)<\ell^{(r)}_{0})+\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{e}(L^{(r)}(0-)\geq\ell^{(r)}_{0})},$ (4.4)
$\alpha^{(r)}_{e}=c^{(r)}_{1}\mu^{(r)}\mathbb{P}(1\leq L^{(r)}\leq\ell^{(r)}_{0})+c^{(r)}_{2}\mu^{(r)}\mathbb{P}(L^{(r)}>\ell^{(r)}_{0}).$ (4.5)
@@
For each integer $k\geq 0$, let $f(z,{\boldsymbol{x}})=z\wedge(k+1)$ for $(z,{\boldsymbol{x}})\in S$, then obviously $f\in C_{b}^{p}(S)$. Since $\Delta_{e}f(X^{(r)}(u)=1(L^{(r)}(u-)\leq k,N^{(r)}_{e}(u)-N^{(r)}_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)}(u)=-1(L^{(r)}(u)\leq k,N^{(r)}_{s}(u)-N^{(r)}_{s}(u-)=1)$, we have
$\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)}(0))\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)}(0))\right]$
@@
Hence, letting $k\to\infty$, we have $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$. Because of this equality, the above formula also yields (4.3). We next let $f(z,{\boldsymbol{x}})=x_{1}$, and apply (4.1), then
$-1+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[T^{(r)}_{e,1}1(L^{(r)}(0-)<\ell^{(r)}_{0})+T^{(r)}_{e,2}1(L^{(r)}(0-)\geq\ell^{(r)}_{0})\right]=0$
@@
Since it follows from (4.4) that
$\min(\lambda^{(r)}_{1},\lambda^{(r)}_{2})\leq\alpha^{(r)}_{e}\leq\max(\lambda^{(r)}_{1},\lambda^{(r)}_{2}),\qquad r\in(0,1],$ (4.7)
@@
$\alpha^{(r)}_{e}$ is uniformly bounded away from 00 and uniformly finite for all $r\in(0,1]$.
To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$, which is possible by Lemma 4.1. We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 5.

### v8

@@ 3.2 Main results
We now present a main result, which will be proved in Section 5. Recall that $L^{(r)}$ is a random variable subject to the stationary distribution of the $r$-th 2-level $GI/G/1$ queue. Define probability distribution $\nu^{(r)}$ under the assumptions (3.a)–(3.e) as
@@
We refer to this $\nu^{(r)}$ as a scaled stationary distribution. Define finite measures $\nu^{(r)}_{i}$ and conditional probability distributions $\nu^{(r)}_{i|cd}$ on $(\mathbb{R}_{+},{\mathcal{B}}(\mathbb{R}_{+}))$ for $i=1,2$ as
$\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{1}),\qquad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;\ell^{(r)}_{1}<L^{(r)}),$
$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{1}),\qquad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|\ell^{(r)}_{1}<L^{(r)}),$
for $B\in{\mathcal{B}}(\mathbb{R}_{+})$. Obviously, $\nu^{(r)}(B)=\nu^{(r)}_{1}(B)+\nu^{(r)}_{2}(B)$.
@@
$h_{1|cd}(x)=1(0\leq x\leq\ell_{1})\times\left\{\begin{array}[]{ll}\displaystyle 1/\ell_{1},&b_{1}=0,\\ \displaystyle\frac{e^{\beta_{1}\ell_{1}}\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}e^{-\beta_{1}x},&b_{1}\not=0,\end{array}\right.$ (3.4)
$h_{2|cd}(x)=1(x>\ell_{1})\beta_{2}e^{-\beta_{2}(x-\ell_{1})}.$ (3.5)
@@
    $\widetilde{A}_{1}=\begin{cases}\displaystyle\frac{2b_{2}\ell_{1}}{c_{2}\sigma_{1}^{2}+2b_{2}\ell_{1}},&b_{1}=0,\\ \displaystyle\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1)},\;&b_{1}\not=0,\end{cases}\qquad\widetilde{A}_{2}=1-\widetilde{A}_{1}.$ (3.10)
@@
Hence, by Little’s law, we have $\rho^{(r)}=\mathbb{P}(L^{(r)}>0)=1-\mathbb{P}(L^{(r)}=0)$. In diffusion applications of a single server queue, the mean queue length is estimated through $(1-\rho^{(r)})\mathbb{E}[L^{(r)}]$ as $r\downarrow 0$. This limit is computed in the following corollary, which is proved in Appendix A.2.
@@
=$\begin{cases}\displaystyle\frac{\sigma_{1}^{2}(2b_{2}\ell_{1}(b_{2}\ell_{1}+c_{2}\sigma_{1}^{2})+c_{2}^{2}\sigma_{1}^{2}\sigma_{2}^{2})}{2(2b_{2}\ell_{1}+c_{2}\sigma_{1}^{2})^{2}},&b_{1}=0,\vspace{1ex}\\ \displaystyle\frac{c_{1}^{2}b_{2}^{2}\sigma_{1}^{2}(e^{\beta_{1}\ell_{1}}-1)+2b_{1}b_{2}\ell_{1}(c_{2}b_{1}-c_{1}b_{2})+c_{2}^{2}b_{1}^{2}\sigma_{2}^{2}}{2b_{1}b_{2}(c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{1}}-1))},\;&b_{1}\not=0.\end{cases}$ (3.15)
@@
$\widetilde{A}_{\exp,1}=\begin{cases}\displaystyle\frac{\beta_{2}\ell_{1}}{1+\beta_{2}\ell_{1}},&b_{1}=0,\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \displaystyle\frac{\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1)},\quad&b_{1}\not=0,\end{cases}$ (3.16)
@@
Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed. However, for a double check, we give its direct proof in Appendix A.3, which is helpful to convince (ii) of Theorem 3.1. This proof is not obvious because the exponential model requires a Markov chain with two-dimensional state space to correctly describe the case that an arrival find the queue length less than $\ell_{1}$ when the previous arrival found the queue length greater than $\ell_{1}$ but multiple departures occurred (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1). This Markov chain will be detailed in Appendix A.3.
@@ 4 BAR’s for the 2-level $GI/G/1$ queue
In this section, we prepare auxiliary notations and results for proving Theorem 3.1. A major difficulty for this proof comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain, and we have to work with $X^{(r)}(\cdot)$. Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions, which will be defined below. This method is called a BAR approach, and has been used in [3, 4, 10]. Recently, this approach is further studied in [12].
Throughout this section, we assume (3.a)–(3.e) unless stated otherwise, but (3.f) is not assumed. The proof of Theorem 3.1 will be done in Section 5.
@@ 4.1 Palm distribution and BAR
Then, for each $r\in(0,1]$, $X^{(r)}(\cdot)$ and the increments of $N^{(r)}_{e}$ and $N^{(r)}_{s}$ are jointly stationary. Since $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite and positive by Lemma 3.1, we can define probability distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ on $(\Omega,{\mathcal{F}})$ by
@@
We now derive a BAR for $X^{(r)}(\cdot)$ using the framework proposed in [12]. For this, we introduce notations for describing the dynamics of $X^{(r)}(\cdot)$. Let $S$ be its state space, that is, $S=\mathbb{Z}_{+}\times\mathbb{R}_{+}^{2}$, and let $C_{b}^{p}(S)$ be the set of all bounded and continuous functions from $S$ to $\mathbb{R}_{+}$ which are partially differentiable from the right concerning the second and third variables and in which $\frac{\partial^{+}}{\partial x_{i}}f$ is the $i$-th partial derivative from the right. Define operator ${\mathcal{H}}$ on $C_{b}^{p}(S)$ which maps $f\in C_{b}^{p}(S)$ to ${\mathcal{H}}f\in C_{b}^{p}(S)$ by
${\mathcal{H}}f({\boldsymbol{x}})=-\frac{\partial^{+}}{\partial x_{2}}f({\boldsymbol{x}})-(c_{1}1(1\leq x_{1}\leq\ell_{1})+c_{2}1(x_{1}>\ell_{1}))\frac{\partial^{+}}{\partial x_{3}}f({\boldsymbol{x}}),\qquad{\boldsymbol{x}}\in S,$
where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})\in\mathbb{R}_{+}^{2}$.
@@
$X^{(r)}_{1}(t)=\begin{cases}X^{(r)}(t-),&\Delta N^{(r)}_{e}(t)=0,\\ (L^{(r)}(t-)+1,T^{(r)}_{e,1}(n_{e,1}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)<\ell^{(r)}_{1},\\ (L^{(r)}(t-)+1,T^{(r)}_{e,2}(n_{e,2}(t)),R^{(r)}_{s}(t-)),\quad&\Delta N^{(r)}_{e}(t)=1,L^{(r)}(t-)\geq\ell^{(r)}_{1},\end{cases}$
@@
where $n_{e,1}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)<\ell^{(r)}_{1})$ and $n_{e,2}(t)=\sum_{n=1}^{N^{(r)}_{e}(t)}1(L^{(r)}(t^{(r)}_{e,n}-)\geq\ell^{(r)}_{1})$. We call $X^{(r)}_{1}(t)$ an intermediate state. Note that
$X^{(r)}(t)=(L^{(r)}_{1}(t)-1,R^{(r)}_{e}(t),T^{(r)}_{s}(N^{(r)}_{s}(t)))$
when $\Delta N^{(r)}_{s}(t)=1$, and $X^{(r)}(t)=X^{(r)}_{1}(t)$ when $\Delta N^{(r)}_{s}(t)=0$.
@@
$\Delta_{e}f(X^{(r)})(t)=f(X^{(r)}_{1}(t))-f(X^{(r)}(t-)),$
$\Delta_{s}f(X^{(r)})(t)=f(X^{(r)}(t))-f(X^{(r)}_{1}(t)).$
Further, let $f(X)(t)=f(X(t))$. Then, from the time evolution of $X^{(r)}(t)$, we have
@@
$+\sum_{u\in(0,t]}[\Delta_{e}f(X^{(r)})(u)+\Delta_{s}f(X^{(r)})(u)],$ (4.2)
where the summation is well defined because $\Delta_{e}f(X^{(r)})(u)$ and $\Delta_{s}f(X^{(r)})(u)$ vanish except for finitely many $u\in(0,t]$. Take the expectation of (4.1) for $t=1$ by $\mathbb{P}$, then the stationarity of $X(\cdot)$ and the definition of the Palm distribution yield
[…]
$\mathbb{E}\left[\sum_{u\in(0,1]}\Delta_{v}f(X^{(r)})(u)]\right]=\alpha^{(r)}_{v}\mathbb{E}^{(r)}_{v}\left[\Delta_{v}f(X^{(r)})(0)\right],\qquad v=e,s.$
@@
$\qquad+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)})(0)\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)})(0)\right]=0,\qquad f\in C_{b}^{p},$ (4.3)
The formula (4.1) is a special case of the rate conservation law (e.g., see [9]), and referred as a basic adjoint relationship, BAR in short, then we call analysis based on this BAR as a BAR approach. A major difficulty in this BAR approach is to evaluate the expectations under the Palm distributions (the last two terms in (4.1)). Their evaluations will be done for the 2-level $GI/G/1$ queue in Section 4.3.
@@
We give two remarks on Palm distributions. (i) $X^{(r)}(0-)$ is not defined under $\mathbb{P}$, but its distribution is well defined under the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{e}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$. So, $X^{(r)}(0-)$ is a random variable which can be defined under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$. Hence, $\Delta_{e}f(X^{(r)})(0)$ and $\Delta_{s}f(X^{(r)})(0)$ in (4.1) are well defined under the Palm distributions.
(ii) Under the Palm distributions, $X^{(r)}_{1}(0)$ is an intermediate state between just after when $N^{(r)}_{e}$ is counted at time 00 and just before when $N^{(r)}_{s}$ is counted at time 00. This means that we take a so called arrival first framework. As for the intermediate state, see [12] for its details. We consider $L^{(r)}(0-)$ under $\mathbb{P}^{(r)}_{e}$, while $L^{(r)}(0)$ under $\mathbb{P}^{(r)}_{s}$. This prevents to directly refer to the intermediate state $X^{(r)}_{1}(0)$ in our arguments. Of course, the intermediate state is not needed if $N^{(r)}_{e}$ and $N^{(r)}_{s}$ do not have a simultaneous count, but such restriction is avoided by introducing $X^{(r)}_{1}(0)$.
@@
Under conditions (3.b) and (3.e),
@@
$\alpha^{(r)}_{e}=\frac{\lambda^{(r)}_{1}\lambda^{(r)}_{2}}{\lambda^{(r)}_{2}\mathbb{P}^{(r)}_{e}(L^{(r)}(0-)<\ell^{(r)}_{1})+\lambda^{(r)}_{1}\mathbb{P}^{(r)}_{e}(L^{(r)}(0-)\geq\ell^{(r)}_{1})},$ (4.5)
$\alpha^{(r)}_{e}=c^{(r)}_{1}\mu^{(r)}\mathbb{P}(1\leq L^{(r)}\leq\ell^{(r)}_{1})+c^{(r)}_{2}\mu^{(r)}\mathbb{P}(L^{(r)}>\ell^{(r)}_{1}).$ (4.6)
@@
For each integer $k\geq 0$, let $f(z,{\boldsymbol{x}})=z\wedge(k+1)$ for $(z,{\boldsymbol{x}})\in S$, then obviously $f\in C_{b}^{p}(S)$. Since $\Delta_{e}f(X^{(r)})(u)=1(L^{(r)}(u-)\leq k,N^{(r)}_{e}(u)-N^{(r)}_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)})(u)=-1(L^{(r)}(u)\leq k,N^{(r)}_{s}(u)-N^{(r)}_{s}(u-)=1)$, we have
$\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[\Delta_{e}f(X^{(r)})(0)\right]+\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[\Delta_{s}f(X^{(r)})(0)\right]$
@@
Since $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$ by Lemma 3.1, the above formula yields (4.4). We next let $f(z,{\boldsymbol{x}})=x_{1}$, and apply (4.1), then
$-1+\alpha^{(r)}_{e}\mathbb{E}^{(r)}_{e}\left[T^{(r)}_{e,1}1(L^{(r)}(0-)<\ell^{(r)}_{1})+T^{(r)}_{e,2}1(L^{(r)}(0-)\geq\ell^{(r)}_{1})\right]=0$
@@
Since it follows from (4.5) and (3.b) that
$0<\lambda^{(r)}_{1}\wedge\lambda^{(r)}_{2}\leq\alpha^{(r)}_{e}\leq\lambda^{(r)}_{1}\vee\lambda^{(r)}_{2}<\infty,\qquad r\in(0,1],$ (4.8)
@@
Under conditions (3.b) and (3.e), $\alpha^{(r)}_{e}$ is uniformly bounded away from 00 and uniformly finite for all $r\in(0,1]$.
To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$ because $\alpha^{(r)}_{s}=\alpha^{(r)}_{e}$ by Lemma 3.1. We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 5.